# LaB-RAG: Label Boosted Retrieval Augmented Generation for Radiology Report Generation

### v1

@@ 3.3 Training LaB-Classifiers for Label Prediction
Because the CheXpert [25] labels are computed using the radiology report, using these labels of the target X-ray in the generation of its corresponding report constitutes a form of data-leakage. We train a set of logistic regression models, LaB-Classifiers, on frozen image embeddings to classify images directly, thereby preventing this leakage (Fig. S1). Further LaB-Classifier training details are in Sec. S3.
@@ 3.4 Label Boosted RAG Algorithm
To generate captions from images, LaB-RAG uses RAG with retrieved example text for ICL. We enhance both retrieval and augmentation steps using categorical labels describing the images and their corresponding text. Given an image at inference time, we rank the similarity of the inference image with all retrieval images in image embedding space (Fig. 2 Top). We apply label-based logic to filter or rerank the similarity scores (Fig. 2 Middle), described in Sec. 3.4.1. We retrieve the corresponding text of the highest ranked images, augment a prompt with the retrieved examples, and autoregressively generate a caption using a pretrained LLM (Fig. 2 Bottom). We additionally incorporate labels in the prompt formatting step (Fig. 2 Middle), described in Sec. 3.4.2. See Alg. 1 for pseudocode.
For RRG, LaB-RAG by default uses BioViL-T [2] projected image embeddings, an “Exact” label filter, the top-5 ranked examples, the “Simple” label format and prompt, and Mistral-7B-Instruct-v0.3 [50]. For generation, we serve the LLM using vLLM [40] and do greedy decoding up to 512 tokens . In all experiments, we use the training and validation splits of our data as our retrieval set and do inference over the test split. Given a retrieval corpus of the target report section, LaB-RAG is able to generate any arbitrary section; however, depending on the radiology report section(s) we aim to generate and compare with, we filter the retrieval and inference sets to only studies with the target section.
@@ 3.4.1 Label Boosted Filtering
LaB-RAG does binary label matching to filter or rerank the ranked list of retrieval samples. LaB-RAG’s label boosting module takes an input list of samples, ranked by image similarity to the inference image, and outputs a ranked list of samples (Alg. 1 Step 8). We present three variations of this module. The simplest variant is “No-filter”, where we do not perform label-based filtering or reordering.
“Exact” filtering requires that every retrieved sample has the same binary labels as the inference image’s labels:
[…]
“Partial” filtering relaxes the constraint of the exact filter by re-sorting the retrieved samples based on the amount of overlapping positive labels between each sample’s label and the inferred image label:
[…]
@@ 3.4.2 Label Boosted Formatting and Prompts
LaB-RAG uses categorical label names as textual image descriptors for image captioning. It does so by formatting the labels as text in the prompt for both the text examples and the inference image. Because the label formatting is intricately associated with the prompt structure, we include the description of our prompts here. In the context of RRG, each CheXpert [25] label is a 14 multiclass multilabel with a 15th binary class label of exclusion (our “Other” label). We thus present four label format and prompt combinations varying the degree of label verbosity and label type description and instruction. The “Naive” prompt does not utilize label descriptors and is closest to a standard RAG prompt with some model instructions.
@@
The “Verbose” prompt additionally includes all label values (negative, uncertain, and unmentioned for CheXpert labels) and expands the model instructions to describe these value types. The prompt does not describe the labels themselves. The “Instruct” prompt uses the exact same template as “Verbose” but adds explicit instructions on how to handle each value type. As the predicted image labels are binary positive or negative, the label set for the inference image will never have labels listed under other values. We refer the reader to Sec. S5 for the examples of the exact prompts and label formats.
Figure 4: Comparison of simple individual label boosting components of LaB-RAG. Compared to SFT methods (Fig. 3), we attain a comparable ${\sim}5\%$ improvement in F1CheXbert using minimal additional computation over standard RAG.
@@ 4 Evaluation Strategy
To evaluate and compare LaB-RAG, we adopt RRG-specific metrics, F1-RadGraph [27] and F1-CheXbert [70]. These measure clinical relevance by computing the F1-score between clinical annotations of the actual and generated reports. The CheXbert annotations are the binarization of the 14 CheXpert [25] labels, while RadGraph annotations are broader in scope. We use the radgraph v0.1.12 and f1chexbert v0.0.2 python packages for their implementations of RadGraph and CheXbert. We also measure natural language metrics, using huggingface evaluate v0.4.2 to compute BLEU-4 [60], ROUGE-L [43], and BERTScore [82]. We refer the reader to our code and Sec. S4 for our specific usage of these metrics.
@@ 4.1 Comparison to Literature Models
We compare against both retrieval-based (CXR-RePaiR [18], CXR-ReDonE [64], X-REM [28]) and direct latent image embedding tuned models (CXRMate [56, 57], CheXagent [12], RGRG [73]). For CXRMate, we specifically use the version submitted to the BioNLP workshop [15, 57]. As each method may have slightly different filtering strategies for data preprocessing, we evaluate on the intersection of the test data splits for each method. LaB-RAG’s data preprocessing only requires there be a ground truth reference text to compare against, thus our test split is a superset of other methods’ test splits. More detailed descriptions of preparing and running each method from the literature are presented in Sec. S6.
As discussed in Sec. 2, RRG models can be categorized by the report section they generate. LaB-RAG is able to generate any target section given a corresponding retrieval corpus and thus comparison to other models only depend on code and data availability. We report results on generated “Findings”, “Impression”, and the concatenation of these sections generated independently. All comparison models were trained in-part over MIMIC-CXR [35].
FactMM-RAG [72] and CXR-RePaiR-Gen [65] are also retrieval based methods, however they do not share code. While BioViL-T [2] and BioViL [4] report RRG metrics and are usable as encoders, they do not share code for autoregressive or precise retrieval based generation. Finally, the only missing component of Pragmatic Retrieval/Llama [54] is code to extract the required indication section. As the MIMIC Code Repository [34] requires modification to extract indications, we are unable to exactly replicate Pragmatic’s precise method.
@@ 5 Studies and Experiments on LaB-RAG
In the following sections, we present results and analyses of our studies on LaB-RAG. These include comparisons of LaB-RAG to literature models and experiments on varying components of our framework. Tab. S10 is a full table of our results across all metrics, models, and experiments. We discuss limitations and next steps where relevant.
@@ 5.1 Baseline Comparisons
We consider models from the literature as our baseline. For models which we could compare against that generate the “Findings” section, none are retrieval based. On “Findings” generation (Fig. 3 Top), LaB-RAG does comparable to SFT methods (RGRG [73], CheXagent [12], CXRMate [57]) on F1CheXbert, however underperforms in F1RadGraph. Yet on “Impression” generation (Fig. 3 Bottom), LaB-RAG is not significantly worse than any SFT method (CheXagent, CXRMate) and outperforms previous retrieval-based methods (CXR-RePaiR [18], CXR-ReDonE [64], X-REM [28]). Across natural language metrics (Fig. S3), differences in performance, while statistically significant, are small in magnitude (Tab. S10).
@@
LaB-RAG In comparison with the previous radiograph, there is little change in the appearance of the heart and lungs. Cardiac silhouette is mildly prominent (cardiomegaly) and there is dense calcification within the aortic arch and descending portion. Blunting of the costophrenic angle is again seen, but no evidence of acute focal pneumonia or vascular congestion. Multiple triple-lumen catheter (supp. dev.) extends to the right atrium.
@@
Table 2: Example “Findings” section demonstrating overvaluation by F1CheXbert. Blue entities are annotated with their corresponding CheXbert label and are also extracted by RadGraph. Red entities are only extracted by RadGraph. Despite CXRMate attaining higher F1RadGraph than LaB-RAG ($0.6$ vs $0.2$) on this example, LaB-RAG matches CXRMate in F1CheXbert (both $1.0$) due to LaB-RAG’s usage of labels.
5.2 Variations of LaB-RAG
We experiment with the core label-boosting components of LaB-RAG and other modular components. Full results across all metrics are available in Sec. S7.
Label Filter and Format Effects. We first consider the simplest strategies for using labels, either “Exact” label match filtering or “Simple” positive label formatting. We show that inclusion of labels marginally changes F1RadGraph results while improving F1CheXbert performance by ${\sim}5\%$ over standard RAG for both “Findings” and “Impression” (Fig. 4). A comparable gain in performance from the literature between CheXagent [12] and CXRMate [57] on “Findings” generation (Fig. 3 Top) required complex new training strategies and additional datasets. As previously observed, LaB-RAG is also comparable to CXRMate in absolute measures. For LaB-RAG, we attain this improvement over standard RAG with minimal additional computation, only requiring “Simple” label formatting; however we do not observe a meaningful improvement when adding “Exact” label filtering to “Simple” formatting, the default LaB-RAG settings. We hypothesize that this is because the LLM is able to infer which parts of the example reports are useful to describe those labels.
To test this, we present inexact label matched examples to the LLM by using our “Partial” label filter (Sec. 3.4.1). Fig. S5 shows that with the default “Simple” prompt, the filters are not meaningfully different in performance. We confirm that inexactly labeled examples are selected by examining the image similarity rank of the filtered selections. In Fig. S10, we observe the “Exact” filter selects fewer than the “Partial” filter of the most visually similar examples. Thus the “Partial” filter will present more mismatched labels to the LLM, yet the stable performance implies the language model is attending to only relevant parts of the examples. Ultimately, given the “Exact” label filter’s improvement over standard RAG and the lack of detriment for the “Simple” label format, we use both label boosting components as defaults for LaB-RAG, aiming to potentially make it simpler for the LLM to parse high-quality label-matched text in the prompt.
Related to prompt complexity, we find that increasing the verbosity of the label format beyond positive labels with the “Simple” prompt tends to decrease performance (Fig. S6).
Image and Language Model Choice. As LaB-RAG is a fully modular framework, we analyze the choice of our pretrained models. We first compare comparable Mistral 7B LLMs: Instruct-v0.1 [29], BioMistral (Instruct-v0.1 further pretrained on PubMed Central) [41], and Instruct-v0.3 [50]. Fig. S7 shows that newer model versions improve RRG performance, while biomedical domain adaptation may be detrimental. This is contrary to previous literature [21] on biomedical domain adaptation, leading us to hypothesize that for a language-intensive approach (ICL), natural language capabilities may be more important than domain specific knowledge. One next step would be to test an LLM with domain-specific instruction following capabilities.
Additionally, the choice of image model directly impacts our label quality. We simulate the effect of using higher accuracy labels by using ground truth labels. We show in Fig. S8 that using true labels improves performance across the board, particularly in impression generation. The gap between LaB-Classifier predicted and true labels (Fig. S2) may be explained by Fig. S10, as image embedding similarity does not necessarily align with the CheXpert [25] labels. A natural extension may be to repeat our experiments with different pretrained image encoders or label sets.
Joint Section Evaluation. Finally, we experiment with modular retrieval examples. Given concatenations of the “Findings” and “Impression” for retrieval, LaB-RAG can jointly generate both sections. We compare this against performance of generating sections independently for studies which have both sections (Fig. S9 Top), showing that LaB-RAG does marginally better when generating jointly. We also compare against CheXagent [12] and CXRMate [57] by evaluating the concatenation of sections each method generates independently (Fig. S9 Bottom). We find a similar result to Fig. S3, where LaB-RAG is comparable to these methods in F1CheXbert but not on other metrics.

### v2

@@ 3.3 Training LaB-Classifiers for Label Prediction
Because the reference labels are extracted from the radiology report, using these labels of the target X-ray in the generation of its corresponding report would constitute data-leakage. We train a set of logistic regression models, LaB-Classifiers, on frozen image embeddings to classify images directly, thereby preventing this leakage (Figure 5). See Section B.2 for further details.
@@ 3.4 Label Boosted RAG Algorithm
To generate captions from images, LaB-RAG uses RAG with retrieved example text for ICL. We enhance both retrieval and augmentation steps using categorical labels describing the images and their corresponding text. Given an image at inference time, we rank the similarity of the inference image to all retrieval images in image embedding space (Figure 2 Top). We apply label-based logic to filter or rerank the similarity scores (Figure 2 Middle), described in Section 3.4.1. We retrieve the corresponding text of the highest ranked images and augment a prompt with the retrieved examples and their labels (Figure 2 Middle), described in Section 3.4.2. The formatted prompt is input to a pretrained, frozen LLM to generate the target caption (Figure 2 Bottom). See Algorithm 1.
For RRG, LaB-RAG by default uses the following modular components: (1) image embeddings from domain and dataset adapted models (BioViL-T for MIMIC-CXR and GLoRIA for CheXpert Plus), (2) inference label predicted by an ensemble of small logistic classifiers, (3) extracted CheXbert retrieval labels, (4) an “Exact” label filter, (5) retrieval of the top-5 ranked examples, (6) the “Simple” label format and prompt, and (7) a general domain generate language model, Mistral-7B-Instruct-v0.3 (MistralAI, 2024). Given a retrieval corpus of the target report section, LaB-RAG is able to generate any arbitrary section; for our experiments, we filter the retrieval and inference sets to only studies with the target section.
@@ 3.4.1 Label Boosted Filtering
LaB-RAG does binary label matching to filter or rerank the ranked list of retrieval samples. LaB-RAG’s label boosting module takes an input list of samples, ranked by image similarity to the inference image, and outputs a ranked list of samples (Algorithm 1 Step 8). We experiment with three variations of this module. The simplest variant is “No-filter”, where we do not perform label-based filtering or reordering.
“Exact” filtering requires that retrieved sample labels match the inference image’s labels:
$\underset{\texttt{exact}}{\texttt{filter}}(\vec{r},\hat{b}_{\mathit{inf}})=\langle s_{x}\in\vec{r}\mid b_{x}=\hat{b}_{\mathit{inf}}\rangle$ (1)
where $\vec{r}$ is a sorted list of samples $s$, $b_{x}$ is the binary label set of a sample $s_{x}$, and $\hat{b}_{\mathit{inf}}$ is inference image’s predicted binary label set. This filtering will most often result in a shorter list than the input $\vec{r}$ and it is possible that the output will be an empty list if the inferred label does not match any labels in the retrieval set (e.g. if the inferred label is unrealistic: both positive “No Finding” and “Atelectasis” in the context of RRG).
“Partial” filtering relaxes the constraint of the exact filter by re-sorting the retrieved samples based on the count of overlapping positive labels between each sample’s label and the inferred image label:
$\begin{gathered}\underset{\texttt{partial}}{\texttt{filter}}(\vec{r},\hat{b}_{\mathit{inf}})=\langle s_{x}\in\vec{r}\mid f(b_{x})\geq f(b_{x+1}),\texttt{idx}_{\,\vec{r}}\,(s_{x})<\texttt{idx}_{\,\vec{r}}\,(s_{x+1})~\text{if}~f(b_{x})=~f(b_{x+1})\rangle\\ f(b)=|b\cap\hat{b}_{\mathit{inf}}|\end{gathered}$ (4)
where $\vec{r}$, $s_{x}$, $b_{x}$, and $\hat{b}_{\mathit{inf}}$ are the same as above, and $f(b)$ is the count of overlapping positive labels between $b$ and $\hat{b}_{\mathit{inf}}$. The second condition gives a stable reordering of $\vec{r}$, such that two samples with the same number of overlapping positive labels retain their relative position from image similarity. Notably, the number of overlapping positive labels does not depend on which labels overlap, nor does it consider the number of overlapping negative labels. This means that two samples with the same number of positive labels overlapped with the image labels can have different positive labels compared to each other and any sample may have more or less total positive labels compared with the image labels. For the “Partial” filter, the lengths of the input and output list are equal.
@@ 3.4.2 Label Boosted Formatting and Prompts
LaB-RAG uses categorical label names as textual image descriptors for image captioning. It does so by formatting the labels as text in the prompt for both the retrieved text examples and the inference image. Because the label formatting is intricately associated with the prompt structure, we include the description of our prompts here. In the context of RRG, each CheXpert/CheXbert-style label is a 14 multiclass multilabel with a 15th binary class label of exclusion (our “Other” label). We thus present four label format and prompt combinations varying the degree of label verbosity and label type description and instruction. The “Naive” prompt does not utilize label descriptors and is closest to a standard RAG prompt with some model instructions.
@@
The “Verbose” prompt additionally includes all label values (negative, uncertain, and unmentioned for CheXpert labels) and expands the model instructions to describe these value types. The prompt does not describe the labels themselves. The “Instruct” prompt uses the exact same template as “Verbose” but adds explicit instructions on how to handle each value type. As the predicted image labels are binary positive or negative, the label set for the inference image will never have labels listed under other values. Section C lists examples of the precise prompts and label formats.
@@ 4 Evaluation Strategy
To evaluate and compare LaB-RAG, we adopt RRG-specific metrics, F1-RadGraph (Jain et al., 2021) and F1-CheXbert (Smit et al., 2020). These measure clinical relevance by computing the F1-score between model derived annotations of clinical terms between the actual and generated reports. The CheXbert annotations are the binarization of the 14 CheXpert/CheXbert labels, while RadGraph annotations are broader in scope. We use the radgraph v0.1.12 and f1chexbert v0.0.2 python packages for their implementations of RadGraph and CheXbert. We also measure natural language metrics, using huggingface evaluate v0.4.2 to compute BLEU-4 (Papineni et al., 2002), ROUGE-L (Lin, 2004), and BERTScore (Zhang* et al., 2020). We refer to our code and Section B.4 for additional details. We report independent results on “Findings” and “Impression” sections.
@@ 4.1 Comparison to Literature Models
We compare against both retrieval-based (CXR-RePaiR (Endo et al., 2021), CXR-ReDonE (Ramesh et al., 2022b), X-REM (Jeong et al., 2024)) and direct latent image embedding tuned models (CXRMate (Nicolson et al., 2024a; b), CheXagent (Chen et al., 2024b), RGRG (Tanida et al., 2023)). For CXRMate, we specifically use the version submitted to the 2024 BioNLP workshop (Demner-Fushman et al., 2024; Nicolson et al., 2024b). As each method may have slightly different filtering strategies for data preprocessing, we evaluate on the intersection of the test data splits for each method. LaB-RAG’s data preprocessing only requires there be a ground truth reference text to compare against, thus our test split tends to be a superset of other methods’ test splits. As not all of these models were developed over the CheXpert Plus (Chambon et al., 2024) dataset, we only compare against performance over MIMIC-CXR (Johnson et al., 2019) which was included in all selected models’ training. More detailed descriptions of preparing and running each method from the literature are presented in Section B.5.
@@ 5 Studies and Experiments on LaB-RAG
Figure 3: Left: LaB-RAG beats other retrieval methods (CXR-RePaiR/ReDonE, X-REM) on RRG metrics. On F1CheXbert, LaB-RAG achieves SOTA on “Findings” generation and performs no different than SFT methods on “Impression” generation (CheXagent, CXRMate). Right: Ablation of individual label boosting components of LaB-RAG. With minimal additional complexity over standard RAG, LaB-RAG has greater gain in F1CheXbert on “Findings” than alternate SFT methods.
In the following sections, we present results and analyses of our studies on LaB-RAG including comparisons of LaB-RAG to literature models and experiments on varying modular components of our framework over both MIMIC-CXR and CheXpert Plus. Tables 12 and 13 show full experimental results across all metrics, with corresponding significance figures in Sections E.1 and E.2.
@@ 5.1 Baseline Comparisons
As a baseline, we compare LaB-RAG to models from the literature over MIMIC-CXR (Figure 3 Left). LaB-RAG achieves state of the art (SOTA) on F1CheXbert on “Findings” generation, compared to SFT methods (RGRG (Tanida et al., 2023), CheXagent (Chen et al., 2024b), CXRMate (Nicolson et al., 2024b)), however underperforms in F1RadGraph. Similarly on “Impression” generation, LaB-RAG does significantly better than CheXagent and comparably to CXRMate on F1CheXbert but worse in F1RadGraph. We do observe that no model performs absolutely well on F1RadGraph for either section, achieving at highest F1RadGraph $0.25$; as established by Yu et al. (2023), this translates to approximately 3 major errors in the report, as determined by board-certified radiologists. Notably, LaB-RAG does significantly better than other retrieval-based methods benchmarked (CXR-RePaiR (Endo et al., 2021), CXR-ReDonE (Ramesh et al., 2022b), X-REM (Jeong et al., 2024)). Differences across natural language metrics are small in magnitude (Figure 7).
Furthermore, we consider the lift of our modular label boosting components over standard RAG (Figure 3 Right). We find that on F1CheXbert, the “Exact” label filter or “Simple” label format both result in a comparable improvement compared with standard RAG, however the effects of the two label boosts are additive, particularly on “Findings” generation. When comparing to literature models on generating the “Findings” section, standard RAG is comparable to CheXagent on F1CheXbert. CXRMate (Nicolson et al., 2024a) achieves a 5.7% gain over CheXagent (Chen et al., 2024b), however this required complex training strategies and additional data. LaB-RAG achieves upwards of an 11% lift over both CheXagent and standard RAG for minimal additional computation. Full results shown in Figures 8 and 15; we note that the small test set size of CheXpert Plus leads to less significant effects with wider standard errors, particularly for “Findings”.
Figure 4: Left: Domain and dataset specificity of image embeddings significantly improves LaB-RAG generations. Right: Improving labeler quality significantly improves LaB-RAG generations. Extr: Extracted from inference target’s ground-truth report, Pred: Predicted from inference image, cxB: CheXbert derived labels, cxP: CheXpert derived labels. For predicted labels, classifiers were trained over labels derived from either the CheXbert or CheXpert labeler.
5.2 Variations of LaB-RAG
Label Filter and Label Format. In Figures 8 and 15, we observe that the “Exact” label filter or the “Simple” label format alone result in comparable results. We note that when using only the “Simple” label format, the labels of the retrieved samples may not be relevant towards the target image. In this setting, we hypothesize that the LLM is able to attribute which parts of the example reports are relevant for the corresponding labels. To test this in the context of a non-trivial label filter and the “Simple” label format, we present inexact label matched examples to the LLM by using our “Partial” label filter. Figures 9 and 16 show that with the default “Simple” prompt, the filters are not meaningfully different in performance. We confirm that inexactly labeled examples are selected by examining the image similarity rank of the filtered selections. In Figure 6, compared to the “Partial” filter, we observe the “Exact” filter selects fewer of the most visually similar examples. In other words, the “Partial” filter presents more mismatched labels to the LLM, yet the stable performance implies the language model is attending to only relevant parts of the examples. Thus small models in the form of the “Exact” label filter and the “Simple” label format focus and contextualize the retrieved examples, synergizing with LLMs with strong natural language capabilities.
Similar to alternate label filters, we find that alternate label formats besides the “Simple” format either result in worse or no different performance (Figures 10 and 17).
Language Model Choice and Number of Retrieved Samples. Motivated by the finding that strong natural language comprehension enables enhanced LaB-RAG performance for RRG, we experiment with alternate language models: Mistral-7B-Instruct-v0.1 (Jiang et al., 2023), BioMistral (Instruct-v0.1 further pretrained on PubMed Central) (Labrak et al., 2024), and Mistral-7B-Instruct-v0.3 (MistralAI, 2024) (the default for our experiments). We find that generally newer model versions improve RRG performance, while biomedical domain adaptation may be detrimental (Figures 11 and 18). This is contrary to previous literature (Gu et al., 2021) on biomedical domain adaptation, leading us to again hypothesize that for a language-intensive approach (RAG and ICL), improved natural language capabilities may be more important than domain specific knowledge (Fan et al., 2024). Future work may include experimenting with an LLM with RRG-specific instruction following capabilities.
Experiments to quantify the effect of retrieving more or less example reports demonstrate minimal differences overall and slight dropoff with fewer than 5 in select settings (Figures 14 and 21).
Image Embedding Quality and Label Quality. As we observe that improvements to the language model enhances generations, we next experiment with image model (Figure 4 Left) and label quality (Figure 4 Right). First, we find that a domain-specific image model (i.e. BioViL-T or GLoRIA) drastically improves embeddings for downstream performance (full results in Figures 12 and 19). Specifically, the embeddings are used for similarity-based retrieval and to train a set of logistic classifiers. The other input for training the classifiers are the “ground-truth”, extracted labels. Following Smit et al. (2020), we observe that CheXbert labels yield better results than CheXpert labels (full results in Figures 13 and 20). We further simulate the effect of using even higher quality labels by using labels extracted from the ground-truth report, providing a theoretical maximum of solely improving classifier performance. We find that directly using extracted labels improves performance across the board, and would result in SOTA performance. Interestingly, even when using CheXbert extracted labels, we do not achieve perfect F1CheXbert; we hypothesize that this small gap may be a result of the LLM ignoring labels presented in the prompt or noise in CheXbert labeling of either the ground-truth or generated reports. Further work is needed to elucidate this observation. Overall, these findings support the claim that LaB-RAG is complementary to SFT methods which may individually or holistically improve the modular components of the method, such as improving the language model, image embeddings, or classification accuracy.
Table 2: Selected “Findings” report section exemplifying differences between F1CheXbert and F1RadGraph. The corresponding CheXbert label is given for blue entities; these entities are also used by RadGraph. Red entities are only extracted by RadGraph. LaB-RAG and CXRMate both score perfect F1CheXbert, but CXRMate attains F1RadGraph $0.63$ vs LaB-RAG’s $0.48$.
@@ 5.2 Variations of LaB-RAG
LaB-RAG The lung volumes are low. This is accentuating the cardiomediastinal silhouette, although there is likely moderate-to-severe cardiomegaly (cardiomegaly). The mediastinum is prominent, which could be due to technique. A right internal jugular catheter (supp. dev.) is present with the tip in the low SVC. There is no pneumothorax. The lungs are clear without consolidation or edema. There is no pleural effusion.